OpenStudy (anonymous):

How are the sin 20° and the cos 70° related?

OpenStudy (aravindg):

sin 20=cos(90-20)=cos 70

OpenStudy (anonymous):

they are equal to each other

OpenStudy (aravindg):

You can see that when they add upto 90 they mean equal values. 20+70=90

OpenStudy (anonymous):

sin and cos are cofunctions, so the sin of one acute angle of a right triangle equals the cos of the other acute angle of a right triangle.